Circular No. 7592

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001Y IN NGC 3362 W. D. Li and M. Modjaz, University of California at Berkeley (UCB),

on behalf of LOTOSS (cf. IAUC 7514), report the discovery of a supernova (mag ∼ 18.1) on unfiltered images taken on Mar. 3.34 and 3.47 UT with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 2001Y is located at α = 10^{h}44^{m}50^{s}.42, δ = +6^{o}35′50′′.8 (equinox 2000.0), which is 18′′.0 west and 2′′.3 north of the nucleus of NGC 3362. SN 2001Y was confirmed at about the same brightness on an unfiltered image taken by M. Schwartz on Mar. 4.2 with the Tenagra Observatory 0.5-m automatic telescope. Images of the same field taken with the Tenagra 0.5-m telescope on Feb. 19.3 (limiting mag ∼ 18.5) and with KAIT on 1999 Apr. 10.3 (limiting mag ∼ 19.5) showed nothing at this position. R. Chornock (UCB), Li, and A. V. Filippenko (UCB) obtained a noisy spectrum of SN 2001Y with the Lick Observatory Shane 3-m telescope on Mar. 3.5 under poor conditions. The spectrum is featureless with a very blue continuum, suggesting that SN 2001Y is a type-II supernova before maximum light.

SUPERNOVA 2001Z IN IC 3528 Modjaz and Li also report the LOTOSS discovery of another apparent

supernova (mag ∼ 18.3) on an unfiltered image taken with KAIT on Mar. 3.5 UT. SN 2001Z is located at α = 12^{h}34^{m}55^{s}.87, δ = +15^{o}34′07′′.5 (equinox 2000.0), which is 0′′.5 west and 11′′.2 north of the nucleus of IC 3528. The new object was confirmed (mag ∼ 18.5) on an unfiltered image taken by Schwartz on Mar. 4.4 with the Tenagra 0.5-m telescope. A KAIT image of the same field taken on Feb. 6.4 showed nothing at the position of the new object (limiting mag ∼ 19.5).

V4334 SAGITTARII C. Lederle and S. Kimeswenger, University of Innsbruck, write: “Near-

infrared photometry of V4334 Sgr obtained on Feb. 18.39 UT at the European Southern Observatory 1-m telescope (+ DENIS) yields magnitudes I > 18, J = 12.61, K = 6.24. This brightening from the deep minimum in summer 2000 (J > 15, K > 7; P. Whitelock, priv. comm.) either indicates an unexpected early opening from the strong circumstellar dust envelope or a renewal of the wind lines of He in the J band. Near-infrared spectroscopic follow-up is strongly encouraged.”

2001 March 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green